OpenStudy (anonymous):

Determine g-1(x) of the function g(x) = 1/2x + 6 where x≠ -3.

OpenStudy (anonymous):

i got 4.5

OpenStudy (anonymous):

if by g-1(x), you mean 'inverse of g' which is g^(-1) (x): switch the positions of y and x from the original equation: y = (1/2)x + 6 switch: x = (1/2)y + 6 solve for y which is also the inverse of g(x) x - 6 = (1/2)y 2x - 3 = y g^(-1)(x) = 2x - 3